# Deza Cayley graphs from difference sets

### v1

@@ 3.2. Constructions from RDSs
    $(252,235,220,218)$
    over the generalized dihedral groups associated with the groups $\mathbb{Z}_{4}^{k-1}\times\mathbb{Z}_{2}$, $\mathbb{Z}_{\frac{q^{2}-1}{2}}$, and $\mathbb{Z}_{126}$, respectively, where $k\geq 3$ and $q\geq 5$ is an odd prime power.
  3. (3)
    There exist a strongly regular Cayley graph with parameters $(16,5,2,0)$ and a strictly Deza Cayley graph with parameters $(16,11,8,6)$ over each of the generalized dihedral groups associated with the groups $\mathbb{Z}_{4}\times\mathbb{Z}_{2}$ and $\mathbb{Z}_{2}^{3}$.
@@
Statement $(1)$ of Proposition 3.5 applied to the RDS from Statement $(1)$ of Lemma 3.6 with $k=3$ leads to a strictly Deza Cayley graph with parameters $(64,11,2,0)$, whereas Statement $(1)$ of Proposition 3.5 applied to the RDSs from Statement $(2)$ of Lemma 3.6 with $q=5$ and $q=9$ leads to non-strictly Deza Cayley graphs with parameters $(24,6,2,0)$ and $(80,12,2,0)$, respectively. Thus, Statement $(1)$ of the corollary holds.
Statement $(2)$ of the corollary is a consequence of Statement $(2)$ of Proposition 3.5 applied to the RDSs from Statements $(1)$$(2)$, and $(4)$ of Lemma 3.6. Finally, Statement $(3)$ of the corollary follows from Statements $(1)$ and $(2)$ of Proposition 3.5 applied to the RDSs from Statement $(3)$ of Lemma 3.6. ∎
The strongly regular Cayley graph from Statement $(3)$ of Corollary 3.7 is isomorphic to the van Lint-Schrijver graph (see [4, p. 371]). All the strictly Deza Cayley graphs from Corollary 3.7 with at most $60$ vertices appear in the list from [11]. We are not able to find other RDSs which satisfy the conditions from Proposition 3.5 and hence can be used for constructing Deza Cayley graphs.
3.3. Two operations
Proposition 3.8.
Let $H$ be a finite group and $\Gamma=\operatorname{Cay}(H,S)$ a strictly Deza Cayley graph over $H$ with parameters $(v,k,2k-v+2,2k-v)$, where $v=|H|$ and $k=|S|$.
  1. (1)
    The graph $\Gamma^{\prime}=\Gamma^{\prime}(m)=\operatorname{Cay}(G,S^{\prime})$, where $G$ is an arbitrary group having $H$ as a subgroup, $m=|G:H|$, and $S^{\prime}=S\cup(G\setminus H)$, is a strictly Deza Cayley graph over $G$ with parameters
    $(mv,k+(m-1)v,(m-2)v+2k+2,(m-2)v+2k).$
  2. (2)
    The graph $\Gamma^{\prime\prime}=\operatorname{Cay}(H\times C,S^{\prime\prime})$, where $C\cong\mathbb{Z}_{2}$, $S^{\prime\prime}=S\cup H^{\#}c$, and $c$ is the nontrivial element of $C$, is a strictly Deza Cayley graph over $H\times C$ with parameters
    $(2v,k+v-1,2k,2k-2).$
Proof.
Let us prove Statement $(1)$. Clearly, $|G|=|G:H||H|=mv$ and $|S^{\prime}|=|S|+|G|-|H|=k+(m-1)v$. So $\Gamma^{\prime}$ is a $(k+(m-1)v)$-regular graph on $mv$ vertices. Due to Lemma 2.1, there is a nonempty proper subset $X$ of $H^{\#}$ such that $S\notin\{X,H^{\#}\setminus X\}$ and
(6) $\underline{S}^{2}=ke+(2k-v+2)\underline{X}+(2k-v)(\underline{H}^{\#}-\underline{X}).$
A straightforward computation in the group ring of $G$ using Eqs. (1) and (6) implies that
$(\underline{S}^{\prime})^{2}=(\underline{S}+\underline{G}-\underline{H})^{2}=\underline{S}^{2}+(\underline{G}-\underline{H})^{2}+2\underline{S}\cdot(\underline{G}-\underline{H})=$
$=ke+(2k-v+2)\underline{X}+(2k-v)(\underline{H}^{\#}-\underline{X})+(m-1)v\underline{H}+(m-2)v(\underline{G}-\underline{H})+2k(\underline{G}-\underline{H})=$
$=(k+(m-1)v)e+((m-2)v+2k+2)\underline{X}+((m-2)v+2k)(\underline{G}^{\#}-\underline{X}).$
The above equality, the condition $S\notin\{X,H^{\#}\setminus X\}$, and Lemma 2.1 yield that $\Gamma^{\prime}$ is a strictly Deza graph with the required parameters.
Now let us prove Statement $(2)$. One can see that $|H\times C|=|H||C|=2v$ and $|S^{\prime\prime}|=|S|+|H|-|1|=k+v-1$ which means that $\Gamma^{\prime\prime}$ is a $(k+v-1)$-regular graph on $2v$ vertices. One can compute using Eqs. (1), (2), and (6) that
$(\underline{S}^{\prime\prime})^{2}=(\underline{S}+\underline{H}^{\#}c)^{2}=\underline{S}^{2}+(\underline{H}^{\#})^{2}+2\underline{S}\cdot\underline{H}^{\#}c=$
$=ke+(2k-v+2)\underline{X}+(2k-v)(\underline{H}^{\#}-\underline{X})+(v-1)e+(v-2)\underline{H}^{\#}+2(k\underline{H}-\underline{S})c=$
$=(k+v-1)e+2k(\underline{X}+\underline{H}c-\underline{S}c)+(2k-2)(\underline{H}^{\#}-\underline{X}+\underline{S}c).$
The latter equality, the condition $S\notin\{X,H^{\#}\setminus X\}$, and Lemma 2.1 imply that $\Gamma^{\prime\prime}$ is a strictly Deza graph with the required parameters. ∎
One can see that the graphs $\Gamma^{\prime}$ and $\Gamma^{\prime\prime}$ again satisfy the condition of Proposition 3.8. It is easy to verify that the graph $\Gamma^{\prime}$ from Statement $(1)$ of Proposition 3.8 is isomorphic to the lexicographic product or composition (see, e.g., [8, p. 398] for the definition) $K_{m}[\Gamma]$ of the clique $K_{m}$ of size $m$ and $\Gamma$. So one can also use [8, Proposition 2.3] to prove Statement $(1)$. The graph $\Gamma^{\prime\prime}$ from Statement $(2)$ of Proposition 3.8 is isomorphic to a disjoint union of two copies of $\Gamma$ in which each vertex from the one copy is adjacent to each vertex from the other one except its own copy.
@@ 3.3. Two operations
Let $\mathcal{K}$ be the class consisting of all Deza Cayley graphs from Corollary 3.3 and Corollary 3.7(2) and the Deza Cayley graph with parameters $(16,11,8,6)$ from Corollary 3.7(3). It is easy to see that each graph from $\mathcal{K}$ satisfies the condition of Proposition 3.8, i.e. has parameters of the form $(v,k,2k-v+2,2k-v)$. Let $\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{K}\mkern-1.5mu}\mkern 1.5mu$ be the closure of $\mathcal{K}$ under the unary operations $\Gamma\mapsto\Gamma^{\prime}(m)$, $m\geq 1$, and $\Gamma\mapsto\Gamma^{\prime\prime}$. Proposition 3.8 implies the following corollary.
Corollary 3.9.
Every graph from $\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{K}\mkern-1.5mu}\mkern 1.5mu$ is a strictly Deza Cayley graph.
Using Deza Cayley graphs from Corollary 3.7(1) and the strongly regular Cayley graph from Corollary 3.7(3), one can produce more non-strictly Deza Cayley graphs. Indeed, let $\mathcal{M}$ be the class consisting of the above graphs. One can see that every graph from $\mathcal{M}$ has parameters of the form $(v,k,2,0)$. Let $\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{M}\mkern-1.5mu}\mkern 1.5mu$ be the closure of $\mathcal{M}$ under the unary operations $\Gamma\mapsto\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu_{n}\times\Gamma$, $n\geq 1$, and the binary operation $(\Gamma_{1},\Gamma_{2})\mapsto\Gamma_{1}\times\Gamma_{2}$, where $\times$ denotes the Cartesian product of graphs and $\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu_{n}$ is a complement to a clique of size $n$. It is easy to verify that every graph from $\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{M}\mkern-1.5mu}\mkern 1.5mu$ is a Cayley graph. Due to [8, Theorem 2.8(iii)-(iv)], every graph from $\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{M}\mkern-1.5mu}\mkern 1.5mu\setminus\mathcal{M}$ is a non-strictly Deza graph with parameters of the form $(v,k,2,0)$. Note that $\mathcal{M}$ contains a strictly Deza graph with parameters $(64,11,2,0)$ from Corollary 3.7(1). Thus, we obtain the following statement.
Corollary 3.10.
Every graph from $\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{M}\mkern-1.5mu}\mkern 1.5mu\setminus\mathcal{M}$ is a non-strictly Deza Cayley graph.
@@
In the end of the paper, we would like to mention that it is possible to consider several other Cayley graphs over generalized dihedral groups arising from DSs and RDSs and not covered by Propositions 3.1 and 3.5. Among them, we could find only two families of strictly Deza Cayley graphs which are exactly divisible design graphs (see [13] for the definition) from [17, Corollary 6.3]. Observe also that the graphs from Propositions 3.1 and 3.5 are not divisible design graphs.

### v2

@@ 5. More Deza graphs
(2v,k+v1,2k,2k2),(2v,k+v-1,2k,2k-2),
respectively.
@@
The graph $K_{m}[\Gamma]$ is a Deza graph with the required parameters by [10, Proposition 2.3] applied to the strongly regular graph $\Gamma_{1}=K_{m}$ and the Deza graph $\Gamma_{2}=\Gamma$.
By the definition, the graph $\Delta=\mkern 1.5mu\overline{\mkern-1.5mu\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu\times K_{2}\mkern-1.5mu}\mkern 1.5mu$ is a (k+v1)(k+v-1)-regular graph on $2v$ vertices. Let $u$ and $v$ be distinct vertices of $\Delta$. If $u$ and $v$ belong to the same copy of $\Gamma$, then they have 2kv+22k-v+2 or 2kv2k-v common neighbors in this copy and v2v-2 common neighbors in the other copy. Therefore they have $2k$ or 2k22k-2 common neighbors. Suppose that $u$ and $v$ belong to distinct copies of $\Gamma$. If the copy of $u$ is adjacent to $v$, then $u$ and $v$ have 2k22k-2 common neighbors, whereas otherwise they have $2k$ common neighbors. Thus, $\Delta$ is a Deza graph with parameters (2v,k+v1,2k,2k2)(2v,k+v-1,2k,2k-2). Since $\Gamma$ is a strictly Deza graph, $\Delta$ so is. ∎
Note that if $\Gamma$ is a Cayley graph, then $K_{m}[\Gamma]$, $m\geq 1$, and $\mkern 1.5mu\overline{\mkern-1.5mu\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu\times K_{2}\mkern-1.5mu}\mkern 1.5mu$ so are. Indeed, let $\Gamma=\operatorname{Cay}(H,S)$ for some group $H$ and an identity-free inverse-closed subset $S$ of $H$. Then
$K_{m}[\Gamma]\cong\operatorname{Cay}(G,S\cup(G\setminus H)),$
where $G$ is any group such that $G\geq H$ and $|G:H|=m$, and
$\mkern 1.5mu\overline{\mkern-1.5mu\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu\times K_{2}\mkern-1.5mu}\mkern 1.5mu\cong\operatorname{Cay}(H\times C,S\cup H^{\#}c),$
where $C\cong\mathbb{Z}_{2}$ and $c$ is the nontrivial element of $C$.
@@
Let $\mathcal{K}$ be the class consisting of all Deza graphs from Theorem 1.1 and the Deza graphs from Theorem 1.2 with a nonzero fourth parameter. It can be verified by a straightforward computation that parameters of every graph from $\mathcal{K}$ satisfy the condition of Lemma 5.1, i.e. they are of the form (v,k,2kv+2,2kv)(v,k,2k-v+2,2k-v). One can see that if $\Gamma$ has parameters of the form (v,k,2kv+2,2kv)(v,k,2k-v+2,2k-v), then $K_{m}[\Gamma]$, $m\geq 1$, and $\mkern 1.5mu\overline{\mkern-1.5mu\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu\times K_{2}\mkern-1.5mu}\mkern 1.5mu$ have parameters of this form. Therefore the closure of $\mathcal{K}$ under the operations
$\Gamma\mapsto K_{m}[\Gamma],~m\geq 1,~\text{and}~\Gamma\mapsto\mkern 1.5mu\overline{\mkern-1.5mu\mkern 1.5mu\overline{\mkern-1.5mu\Gamma\mkern-1.5mu}\mkern 1.5mu\times K_{2}\mkern-1.5mu}\mkern 1.5mu$
consists of strictly Deza Cayley graphs.
Let $\mathcal{M}$ be the class consisting of all graphs from Theorem 1.2 with a zero fourth parameter. One can see that every graph from $\mathcal{M}$ has parameters of the form $(v,k,2,0)$. Due to [10, Theorem 2.8(iii)-(iv)], the closure of $\mathcal{M}$ under the operations
$\Gamma\mapsto\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu_{m}\times\Gamma,~m\geq 1~\text{and}~(\Gamma_{1},\Gamma_{2})\mapsto\Gamma_{1}\times\Gamma_{2}$
consists of Deza Cayley graphs. One can verify straightforwardly using the definitions of the above operations that all of these graphs are non-strictly Deza graphs except for the graph with parameters $(64,11,2,0)$ from Theorem 1.2.
6. Concluding remarks
We do not know whether there exist other DSs with parameters satisfying the condition $k=\frac{2v-1-\sqrt{8v-7}}{2}$ which can be used for the construction from Proposition 4.1 and are not covered by Lemma 2.1. Checking using the database of DSs [12] implies that there is no other such a DS with $v\leq 100000$. We are not able to find RDSs with parameters satisfying one of the conditions $(n,\lambda)=(2,2)$, $(n,\lambda)=(4,2)$, $\lambda=2$ which can be used for the construction from Proposition 4.2 and are not covered by Lemma 2.2.
@@ 6. Concluding remarks
We would like to mention that it is possible to consider several other Cayley graphs over generalized dihedral groups arising from DSs and RDSs and not covered by Propositions 4.1 and 4.2. One can obtain necessary and sufficient conditions for these graphs to be Deza graphs in terms of parameters of DSs or RDSs. We could find appropriate RDSs only for two infinite families among these graphs which are exactly divisible design graphs (see [18] for the definition) from [26, Corollary 6.3]. Observe also that the graphs from Propositions 4.1 and 4.2 are not divisible design graphs.